if x=(p+q)/(p-q), then x-1=?
1
Expert's answer
2012-07-17T07:19:35-0400
x-1=(p+q)/(p-q)-1=(p+q-(p-q))/(p-q)=(p+q-p+q)/(p-q)=(q+q)/(p-q)=2q/(p-q)